# Alternative Approaches for Computing Highest-Density Regions

### v1

@@ 4.4.1 Simplex scenario: Dirichlet
A particular interest is devoted to the simplex scenario. We consider a Dirichlet random vector $X\sim\mathcal{D}(\alpha_{1},\alpha_{2},\alpha_{3})$ defined on the $2$-dimensional simplex, where $\alpha_{j}>0,\ j=1,2,3$; we refer to Chapter XI in Devroye, (1986) for a detailed exposition. Up to a normalizing constant, the density of $X$ is given by
@@
with $u_{1},u_{2}$ such that $u_{1},u_{2}\in[0,1]$ and $(1-u_{1})^{\frac{1}{a+1}}+(1-u_{2})^{\frac{1}{a+1}}\geq 1$. The analytical derivation is given in Supplementary Material A, while its graphical representation–in terms of a set of random draws and different level sets–can be depicted in Figure 5.
Figure 5: Contours of a Dirichlet(1,1,2) density (left) and of its copula density (right). The $95\%$ HDR is depicted in blue color.
To evaluate the proposed measures, we first need to check or perform an accurate tuning of their hyperparameters. As shown in the Supplementary Material B, the optimal choice of $k$–which represents the optimal number of neighbors to take into account for computing both $M_{1}$:$k$NN-Eucl and $M_{2}$:$k$NN-CDF–remains the same as for the other scenarios: $k=\left[\sqrt{n/2}\right]$. However, for measures depending on $\epsilon$–which defines the length, area, or volume of the neighborhood–the optimal choice depends on the support of the underlying variable. Being defined on a simplex, in a Dirichlet scenario, the optimal values of $\epsilon$ have a smaller magnitude compared to noncompact or less restrictive scenarios such as S1-S16. In this case, the following choices are made for $\epsilon$, guided by simulation studies reported in the Supplementary Material B:
@@
Performances are robust to sample size, with optimal $\epsilon=0.10$.
@@
Performances depend on the sample size, according to a heuristic given by: $\epsilon=\exp\left(-1.22-0.23\log(n)\right)$. This relationship is obtained by fitting a nonlinear regression (exponential decay model), with the empirical optimal $\epsilon$ and the sample size $n$ as dependent and independent variables, respectively.
@@
Performances are robust to sample size, with optimal $\epsilon=0.02$.
In terms of results, as suggested in Table 1, the more complex the scenario, the more difficult we expect it to be for identifying the highest-density points (or true negatives) vs. the true positives. In this particular case, as shown in Figure 6, the discrepancy between the different measures is even more evident with high and persistently high FNR values, even with increasing sample sizes, for $M_{0}$:KDE and $M_{2}$:$k$NN-CDF. All other measures improve with the sample size, achieving results comparable to the other scenarios.
Figure 6: Performance results in terms of mean and error bounds of the compared methods in the Dirichlet scenario for varying sample sizes.
@@ 4.5 MAGIC data
We now apply the proposed measures to construct an HDR for the joint distribution of two selected variables from the MAGIC dataset (Bock et al.,, 2004). These data simulate the registration of high-energy gamma particles in a ground-based atmospheric Cherenkov gamma telescope, and have been studied in classification problems (Dvořák and Savický,, 2007), as well as to analyze the dependency structure of some of the variables (see e.g., Nagler and Czado,, 2016; Grazian et al.,, 2022).
In this evaluation, we focus on gamma-ray observations (overall $n=12,332$) and consider the two variables “fConc1” and “fM3Long”, after scaling them. We refer to Bock et al., (2004) for a full description of the dataset. In this case (as deduced from the complex structure of the data; see Figure 7), the parametric approach is inappropriate for both the estimation of the marginal distribution and, more importantly, the copula model. Thus, in Figure 7, we illustrate the derived 95% HDR using the nonparametric measures only. Although in the absence of the underlying truth it is not possible to perform a reliable evaluation, it seems that the two nonparametric copula-based approaches ($M^{\text{NPCop}}_{0}$:DE, $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF), as well as the distance-based $M_{1}$:$k$NN-Eucl and $M_{3}$:$\epsilon$-CDF, more sensibly exclude tail data points (which may be expected to have a lower density) from the HDR.
@@
To deal with estimation and data uncertainty, we also explore the potential of a “measure averaging” in a similar, but simplified, fashion to model averaging (see e.g., Hoeting et al.,, 1999; Hjort and Claeskens,, 2003). By forming a consensus between the different available measures, we expect the resulting HDR to be more robust than the individual estimates. Specifically, in Figure 8, we illustrate a 95% HDR formed by averaging across the different nonparametric measures: the estimated HDR is defined by the set of points identified as highest-density points by more than half of the evaluated metrics. Notably, when comparing the resulted HDR in Figure 8 and those obtained by the individual measures (Figure 7), we identify $M_{3}$:$\epsilon$-CDF and $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF as the closest regions (both with < 1% classification difference). Therefore, these may result in a possible superior ability to detect anomalous values.
@@ 5 Discussion and Conclusion
In this work, we have discussed some generalizations of the standard density-based approach to compute HDRs by using neighborhood measures. Various measures, with distinct properties, have been introduced and compared to the classical kernel density estimator, across different scenarios and sample sizes. Furthermore, the use of copula models for representing a multivariate distribution through a combination of its marginals and their dependence structure has been investigated. Our results suggest that such a generalized approach may provide great advantages to HDR estimation when considering several alternative measures to KDE, in particular those based on copulae (both under a parametric and nonparametric approach). In fact, compared to traditional KDE, the copula-based HDRs resulted in greater accuracy and lower FPR and FNR in a number of simulation scenarios and possibly in real data (as suggested by the MAGIC data application). Such performances are particularly important when the interest is in balancing different types of errors, minimizing both false positives and false negatives, and maximizing therefore the probability of detecting outliers or anomalous values. Among the different evaluated measures, those based on the $\epsilon$-neighborhood multivariate CDF distance ($M_{3}$:$\epsilon$-CDF, $M^{\text{PCop}}_{3}$:$\epsilon$-CDF, $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF) show the most promising results, especially in the nonparametric setting and when the copula is involved. Since the main characteristic of these measures is that they are based on the CDF rather than the PDF, it may suggest that focusing on estimating the former may be advantageous. Similar findings are reported by Magdon-Ismail and Atiya, (2002), who acknowledge that, compared to directly estimating the PDF with e.g., KDE, approximating the CDF is less sensitive to statistical fluctuations and its convergence rate is faster than the convergence rate of KDE methods. Although heuristic considerations for each specific measure are made in terms of their hyperparameters, e.g., the choice of $k$ in the kNN-based approaches, further work may investigate the existence of optimal theoretical values in a similar fashion to the plethora of asymptotic works on the optimal bandwidth choice in KDE (see e.g., Wand and Jones, 1994b, ; Chacón et al.,, 2011).
@@
In this work, we focused on estimating HDR for continuous distributions that are dominated by the Lebesgue measure. Furthermore, despite providing a general multidimensional framework for building HDRs, we evaluated the proposed approach in a bivariate context. Future lines of research may examine the problem in the underappreciated setting of discrete probability distributions and in higher dimensions. In the first case, possible connections could be made with the work of O’Neill, (2022), which established some theory and an algorithm for HDRs in the space of discrete distributions. In the second case, and more specifically with reference to copula-based measures, our aim is to explore the use of vine copulae (Nagler and Czado,, 2016) to construct flexible dependence models for an arbitrary number of variables using only bivariate building blocks. We expect, in fact, to see remarkable advantages in using copulae over an increased number of variables, as the extension of the common KDE to high dimensions has proven challenging in terms of both computational efficiency and statistical inference.
Supporting information
Additional supporting information is provided in the Supplementary material, made available at the end of this document. The implementation of the algorithms and R codes to reproduce all the results and Figures/Tables of the main manuscript are provided at https://github.com/nina-DL/HDR_neigh_measures.
@@ Appendix B Supporting Details for Hyperparameter Choices
In this section, we illustrate the choices we made for the hyperparameters of the different measures evaluated in this paper, which we reiterate below. We support our choices with simulation studies in which we compare the performances of our measures for different values of the hyperparameters defined on an appropriate grid.
@@
Cumulative Euclidean distances from the $k$NNs. We sum the Euclidean distances between each point and its $k$ neighbors defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, leading to a general rule of thumb aligned with the existing literature: $k=\left[\sqrt{n/2}\right]$, with $\left[x\right]$ the integer closest to $x$. The supporting simulation results are given in Figures 9-12.
@@
Cumulative CDF distances from the $k$NNs. We sum the CDF distances between each point and its $k$ neighbors, which are defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, suggesting a uniform choice across different scenarios and sample sizes: $k=30$. The supporting simulation results are given in Figures 13-16.
@@
This represents the $\epsilon$-neighborhood multivariate CDF distance, which uses empirical CDF as the CDF estimator. The optimal choice of the hyperparameter $\epsilon$ is based on an extensive cross-validation procedure, leading to the following heuristic for S1-S16 as a result of an exponential decay model fit with respect to the sample size: $\epsilon=\exp\left(2.13-0.3\log(n)\right)$. For S1-S16, the supporting simulation results are given in Figures 17-20. For S17 (Dirichlet scenario), performances are robust to sample size, with optimal $\epsilon=0.10$; see Figure 21.
@@
Fully nonparametric indirect estimation of the $\epsilon$-CDF measure with copula. We use the empirical CDF for estimating the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. The optimal choice of the hyperparameter $\epsilon$ follows the same strategy as the previous measure, leading to the following heuristic for S1-S16: $\epsilon=\exp\left(1.74-0.26\log(n)\right)$. For S1-S16, the supporting simulation results are given in Figures 22-25. For S17 (Dirichlet scenario), performances depend on the sample size, according to a heuristic given by: $\epsilon=\exp\left(-1.22-0.23\log(n)\right)$. This relationship is obtained by fitting a nonlinear regression (exponential decay model), with the empirical optimal $\epsilon$ and the sample size $n$ as dependent and independent variables, respectively; see Figure 26.
@@
Parametric indirect CDF estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. The optimal choice of the hyperparameter $\epsilon$ follows the heuristic $\epsilon=\exp\left(1.60-0.41\log(n)\right)$ for S1-S16 (see Figures 27-30). For S17 (Dirichlet scenario), performances are robust to sample size, with optimal $\epsilon=0.02$; see Figure 31.
@@
Figure 9: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=5$, with $n=50$.
@@
Figure 10: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=7$, with $n=100$.
@@
Figure 11: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=16$, with $n=500$.
@@
Figure 12: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=22$, with $n=1000$.
@@
Figure 13: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 14: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 15: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 16: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 17: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=2.60$.
@@
Figure 18: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=2.11$.
@@
Figure 19: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=1.30$.
@@
Figure 20: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=1.06$.
@@
Figure 21: Performance results in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the optimal choice given by $\epsilon=0.10$.
@@
Figure 22: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)\approx 2$.

### v2

@@ 4.4.1 Simplex scenario: Dirichlet
A particular interest is dedicated to the simplex scenario, which plays an important role as the sample space of compositional data. Compositional data quantitatively describe parts of some whole and consist of vectors of positive components subject to a unit-sum constraint (Aitchison,, 1982). Measurements involving proportions, probabilities, or percentages can all be thought of as compositional data. These commonly arise in many disciplines; for example, in demography, cause-specific mortality rates can be studied by considering them as compositions, where no single rate is free to vary separately from the rest of the rate composition (see e.g., Stefanucci and Mazzuco,, 2022). This induces a particular and unique dependence structure (represented, e.g., by the correspondent copula), which is fully entangled in the whole system.
The Dirichlet distribution represents a natural candidate for analyzing compositional data, since its support is the simplex. Let $X\sim\mathcal{D}(\alpha_{1},\alpha_{2},\alpha_{3})$ denote a Dirichlet random vector defined on the $2$-dimensional simplex, where $\alpha_{j}>0$, for $j=1,2,3$; we refer to Chapter XI in Devroye, (1986) for a detailed exposition. Up to a normalizing constant, the density of $X$ is given by
@@
with $u_{1},u_{2}$ such that $u_{1},u_{2}\in[0,1]$ and $(1-u_{1})^{\frac{1}{a+1}}+(1-u_{2})^{\frac{1}{a+1}}\geq 1$. The analytical derivation is given in Supplementary Material A, while its graphical representation–in terms of a set of random draws and different level sets–is illustrated in Figure 4.
Figure 4: Contours of a Dirichlet(1,1,2) density (left) and of its induced copula density (right). The $95\%$ HDR is represented by the blue contour.
To evaluate the proposed measures in this specific scenario, one first needs to perform an accurate tuning of their hyperparameters. As shown in the Supplementary Material B, the optimal choice of $k$–which represents the optimal number of neighbors to take into account for computing both $M_{1}$:$k$NN-Eucl and $M_{2}$:$k$NN-CDF–remains the same as for the other scenarios: $k=\left[\sqrt{n/2}\right]$. However, for measures depending on $\epsilon$–which defines the length, area, or volume of the neighborhood–the optimal choice depends on the support of the underlying variable. Being defined on a simplex, in a Dirichlet scenario, the optimal values of $\epsilon$ have a smaller magnitude compared to noncompact or less restrictive scenarios such as S1-S16. In this case, the following choices are made for $\epsilon$, guided by simulation studies reported in the Supplementary Material B:
@@
Performances are robust to sample size, with the empirical optimal $\epsilon=0.10$.
@@
Performances depend on the sample size, according to a heuristic given by: $\epsilon=\exp\left(-1.22-0.23\log{n}\right)$. This relationship is obtained by fitting a nonlinear regression (exponential decay model), with the empirical optimal $\epsilon$ and the sample size $n$ as dependent and independent variables, respectively.
@@
Performances are robust to sample size, with the empirical optimal $\epsilon=0.02$.
In terms of results, as suggested in Table 1, the more complex the scenario, the more difficult one should expect it to be to identify the highest-density points (or true negatives) versus the true positives. In this particular case, as shown in Figure 5, the discrepancy between the different measures is remarkable, with persistently high FNR values, even when the sample size increases, for $M_{0}$:KDE and $M_{2}$:$k$NN-CDF. All other measures improve with the sample size, achieving results comparable to the other scenarios.
Figure 5: Performance results of the compared methods in the Dirichlet scenario for varying sample sizes. Data are summarized in terms of mean and error bounds across $1000$ independent Monte Carlo (MC) replicates.
@@ 4.5 MAGIC data
We now apply the proposed measures to derive an HDR for the joint distribution of two selected variables from the MAGIC dataset (Bock et al.,, 2004). These data simulate the registration of high-energy gamma particles in a ground-based atmospheric Cherenkov gamma telescope, and have been studied in classification problems (Dvořák and Savický,, 2007), as well as to analyze the dependence structure of some of the characterizing variables (see e.g., Nagler and Czado,, 2016; Grazian et al.,, 2022).
In this evaluation, we focus on gamma-ray observations (overall $n=12,332$) and consider the two variables “fConc1” and “fM3Long”, after scaling them. We refer to Bock et al., (2004) for a full description of the dataset. In this case (as deduced from the complex structure of the data; see Figure 6), the parametric approach is inappropriate for both the estimation of the marginal distribution and, more importantly, the copula model. Thus the 95% HDR is estimated using the nonparametric measures only. Although in the absence of the underlying truth it is not possible to perform a reliable evaluation, it seems that the two nonparametric copula-based approaches ($M^{\text{NPCop}}_{0}$:DE and $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF), as well as the distance-based $M_{1}$:$k$NN-Eucl and $M_{3}$:$\epsilon$-CDF, more sensibly exclude tail data points (which may be expected to have a lower density) from the HDR.
@@
To deal with estimation and data uncertainty, we also explore the potential of a “measure averaging” in a similar, but simplified, manner to model averaging (see e.g., Hoeting et al.,, 1999; Hjort and Claeskens,, 2003). By forming a consensus between the different available measures, one would expect the resulting averaged HDR to be more robust than the individual estimates. Specifically, in Figure 7, we illustrate a 95% HDR formed by averaging across the different nonparametric measures: the estimated HDR is defined by the set of points identified as highest-density points by more than half of the evaluated metrics. Notably, when comparing the resulting HDR in Figure 7 and those obtained by the individual measures (Figure 6), we identify $M_{3}$:$\epsilon$-CDF and $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF as the regions better resembling the average (both with ¡ 1% classification difference). Therefore, these may result in a possible superior ability to detect anomalous values in scenarios where a parametric assumption is unrealistic.
@@ 5 Discussion and conclusions
In this work, we have discussed some generalizations of the standard density-based approach for estimating highest-density regions, using neighborhood measures. Various measures, with distinct properties, have been introduced and compared to the classical kernel density estimator, across different scenarios and sample sizes. Furthermore, the use of copula models for representing a multivariate distribution through a combination of its marginals and their dependence structure has been investigated. Our results suggest that such a generalized approach may provide great advantages to HDR estimation when considering several alternative measures to KDE, especially those based on copulae. In fact, compared to traditional KDE, copula-based HDR resulted in greater accuracy and lower FPR and FNR in a number of simulation scenarios and possibly in real data (as suggested by the MAGIC data application). Such performances are particularly important when the interest is in balancing different types of errors, minimizing both false positives and false negatives, and maximizing therefore the probability of detecting atypical or anomalous values. In a doping detection problem, for example, this would directly translate into an enhanced ability of identifying doping abuse.
It is important to emphasize that among the various copula-based measures, the parametric ones generally outperformed the non-parametric measures. This outcome was expected because all the considered scenarios, despite their complexity, were generated from parametric models and were evaluated under the assumption of no model misspecification for the marginals. In real-world applications, one would first need to select the most appropriate family using a criterion of fit, such as the Akaike Information Criterion (AIC) or the Bayesian Information Criterion (BIC). Notably, an important advantage of using copulae is that marginal model selection is implemented separately for each marginal, restricting the analysis to simpler univariate cases. Concerning the copula model, we account for potential misspecification, with model selection performed using the AIC (note that built-in functions are available in R software, e.g., the BiCopSelect() function from the VineCopula package (Nagler et al.,, 2023)).
Relative to the scenarios and measures considered in this work, the main recommendation for the neighborhood measure is thus a copula-based alternative. In particular, parametric variants should be preferred when distributions show a pattern traceable to a parametric family. When parametric measures do not appear appropriate for the data at hand, we would recommend the nonparametric copula-based options, with the exception of scenarios with heavy-tailed marginals. Here, $M_{1}$:$k$NN-Eucl achieves the best results, and the nonparametric $M^{\text{NPCop}}_{0}$:DE measure is among the most inferior ones, especially in small samples. On the contrary, the $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF measure using the $\epsilon$-neighborhood multivariate CDF distance shows promising results. Since the main characteristic of the latter is that it is based on the CDF rather than the PDF, it may suggest that focusing on estimating the former may be advantageous. Similar findings are reported by Magdon-Ismail and Atiya, (2002), who acknowledge that, compared to directly estimating the PDF with e.g., KDE, approximating the CDF is less sensitive to statistical fluctuations and its convergence rate is faster than the convergence rate of KDE methods. We also emphasize that although heuristic considerations for each specific measure are made in terms of their hyperparameters, e.g., the choice of $k$ in the kNN-based approaches, further work may investigate the existence of optimal theoretical values in a similar fashion to the asymptotic works on the optimal bandwidth choice in KDE (see e.g., Wand and Jones, 1994b, ; Chacón et al.,, 2011).
@@
In this work, we focused on estimating HDR for continuous distributions that are dominated by the Lebesgue measure. Furthermore, despite providing a general multidimensional framework for building HDRs, we evaluated the proposed approach in a bivariate context. Future lines of research may examine the problem in the underappreciated setting of discrete probability distributions and in higher dimensions. In the first case, possible connections could be made with the work of O’Neill, (2022), which established some theory and an algorithm for HDRs for discrete distributions. In the second case, and more specifically with reference to copula-based measures, our aim is to explore the use of vine copulae (Nagler and Czado,, 2016) to construct flexible dependence models for an arbitrary number of variables using only bivariate building blocks. We expect, in fact, to see remarkable advantages in using copulae over an increased number of variables, as the extension of the common KDE to high dimensions has proven challenging in terms of both computational efficiency and statistical inference.
@@
Additional supporting information is provided in this Supplementary material. The implementation of the algorithms is possible with the R package HDR2D provided in the Github repository https://github.com/nina-DL/HDR2D.
@@ Appendix B Supporting Details for Hyperparameter Choices
In this section, we illustrate the choices we made for the hyperparameters $k$ and $\epsilon$ of the different measures evaluated in this paper, which we reiterate below. We support our choices with simulation studies in which we compare the performances of our measures for different values of the hyperparameters defined on an appropriate grid. The grid is defined on a sequence of generally equi-spaced points and takes into account the admissible choices of the hyperparameters $k$ and $\epsilon$ (that is, $k\in\{1,2,\dots,n\}$ and $\epsilon\geq 0$) and the range of variation as well as the trend induced on the different performance metrics for different sample sizes. For example, in Figure 8, the grid for $k$ when $n=50$ is defined on all integers from $1$ to $50$; in Figure 9 with $n=100$, we consider the same grid, as no improvements are expected for $k>50$ given the observed trend. We let the grid vary with $n$ and the measure, but not with the different scenarios S1-S17. The only exception is made for $\epsilon$ in S17, since Dirichlet’s support is limited to the simplex, and as such must be $\epsilon$. All the grids of reference are reported in the caption of the corresponding figure. All results for each hyperparameter choice are summarized in terms of their mean across $50$ independent Monte Carlo (MC) replicates.
@@
Cumulative Euclidean distances from the $k$NNs. We sum the Euclidean distances between each point and its $k$ neighbors defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, leading to a general rule of thumb aligned with the existing literature: $k=\left[\sqrt{n/2}\right]$, with $\left[x\right]$ the integer closest to $x$. The supporting simulation results are given in Figures 8-11.
@@
Cumulative CDF distances from the $k$NNs. We sum the CDF distances between each point and its $k$ neighbors, which are defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, suggesting a uniform choice across different scenarios and sample sizes: $k=30$. The supporting simulation results are given in Figures 12-15.
@@
This represents the $\epsilon$-neighborhood multivariate CDF distance, which uses empirical CDF as the CDF estimator. The optimal choice of the hyperparameter $\epsilon$ is based on an extensive cross-validation procedure, leading to the following heuristic for S1-S16 as a result of an exponential decay model fit with respect to the sample size: $\epsilon=\exp\left(2.13-0.3\log{n}\right)$. For S1-S16, the supporting simulation results are given in Figures 16-19. For S17 (Dirichlet scenario), performances are robust to sample size, with optimal $\epsilon=0.10$; see Figure 20.
@@
Fully nonparametric indirect estimation of the $\epsilon$-CDF measure with copula. We use the empirical CDF for estimating the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. The optimal choice of the hyperparameter $\epsilon$ follows the same strategy as the previous measure, leading to the following heuristic for S1-S16: $\epsilon=\exp\left(1.74-0.26\log{n}\right)$. For S1-S16, the supporting simulation results are given in Figures 21-24. For S17 (Dirichlet scenario), performances depend on the sample size, according to a heuristic given by: $\epsilon=\exp\left(-1.22-0.23\log{n}\right)$. This relationship is obtained by fitting a nonlinear regression (exponential decay model), with the empirical optimal $\epsilon$ and the sample size $n$ as dependent and independent variables, respectively; see Figure 25.
@@
Parametric indirect CDF estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. The optimal choice of the hyperparameter $\epsilon$ follows the heuristic $\epsilon=\exp\left(1.60-0.41\log{n}\right)$ for S1-S16 (see Figures 26-29). For S17 (Dirichlet scenario), performances are robust to sample size, with optimal $\epsilon=0.02$; see Figure 30.
@@
Figure 8: Performance results of $M_{1}$:$k$NN-Eucl in all evaluated scenarios for different values of the hyperparameter $k$, with $k\in\{1,2,\dots,50\}$. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=5$, with $n=50$.
@@
Figure 9: Performance results of $M_{1}$:$k$NN-Eucl in all evaluated scenarios for different values of the hyperparameter $k$, with $k\in\{1,2,\dots,50\}$. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=7$, with $n=100$.
@@
Figure 10: Performance results of $M_{1}$:$k$NN-Eucl in all evaluated scenarios for different values of the hyperparameter $k$, with $k\in\{1,2,\dots,50\}$. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=16$, with $n=500$.
@@
Figure 11: Performance results of $M_{1}$:$k$NN-Eucl in all evaluated scenarios for different values of the hyperparameter $k$, with $k\in\{1,2,4,6,\dots,98,100\}$. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=22$, with $n=1000$.
@@
Figure 12: Performance results of $M_{2}$:$k$NN-CDF in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $k$, with $k\in\{1,2,4,6\dots,28,30\}$. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 13: Performance results of $M_{2}$:$k$NN-CDF in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $k$, with $k\in\{1,2,4,6\dots,48,50\}$. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 14: Performance results of $M_{2}$:$k$NN-CDF in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $k$, with $k\in\{1,2,4,6\dots,48,50\}$. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 15: Performance results of $M_{2}$:$k$NN-CDF in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $k$, with $k\in\{1,5,10,15\dots,145,150\}$. Gray vertical line refers to the final choice: $k=30$.
@@
Figure 16: Performance results of $M_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log{n}\right)=2.60$.
@@
Figure 17: Performance results of $M_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log{n}\right)=2.11$.
@@
Figure 18: Performance results of $M_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log{n}\right)=1.30$.
@@
Figure 19: Performance results of $M_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log{n}\right)=1.06$.
@@
Figure 20: Performance results of $M_{3}$:$\epsilon$-CDF in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.001,0.010,0.020,\dots,0.290,0.300\}$. Gray vertical line reflects the optimal choice given by $\epsilon=0.10$.
@@
Figure 21: Performance results of $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log{n}\right)\approx 2$.